OpenStudy (anonymous):

What is the area of a triangle with a base of 6.8 cm and a height of 5.3 cm? A. 24.2 cm2 B. 36.04 cm2 C. 18.02 cm2 D. 12.1 cm2

OpenStudy (mathmale):

I'd be happy to help, and am sure others would be too. But first, would you please look up and type in the formula for the area of a triangle of base b and height h?

OpenStudy (jdoe0001):

\(\bf \textit{area of a triangle}=\cfrac{1}{2}\cdot base\cdot height\)

OpenStudy (cggurumanjunath):

A = 1/2 *base*height

OpenStudy (anonymous):

ok...............

OpenStudy (mathmale):

Love to help, but I do expect you to be an active participant in solving this problem.

OpenStudy (anonymous):

really? I NEED help. i tried to learn but got a 20% on my last test an i have to finish 10 lessons before 2:00

OpenStudy (mathmale):

You're on your own, because two people have given you the correct formula. See you again: next time, let's work together to solve the problem you've posted.

OpenStudy (mathmale):

Involving yourself in problem solutions will greatly increase your ability to solve problems. Sorry you don't think I "helped."

OpenStudy (jdoe0001):

your "help" definition may be at odds with most folks "help" definition "help" would stand for, you may not know and thus you'd another pair of eyes or feed as opposed to "do it for me "

OpenStudy (mathmale):

I "helped" by asking our friend SnowFlake to find and type in the appropriate formula; SnowFlake responded with "I'm too lazy." :(

OpenStudy (mathmale):

But thanks for your input, JDoe. We both mean well.

OpenStudy (jdoe0001):

The beginning is the most important part of any work. ~~ Plato (370 BC) ~~

OpenStudy (anonymous):

just leave me alone if you cant help. i am sorry and i wish to be friends. i agree i just wanted to not be the only one who will watch a movie i have been waiting 4. I am sorry can we set this up and be friends? and thank you for your help

OpenStudy (mathmale):

Your being too lazy is anything but a reason for anyone to do your work for you. I really, truly, don't mean to be unkind. Later, when you're older, perhaps the point I'm trying to make will be clearer for you, as well as less offensive. You could have found the formula simply by doing a 'Net search for "area of a triangle." How long would that take (compared to your expressing your anger at me because I'm not "helping" you in the way you wanted?

OpenStudy (anonymous):

i am sorry but can you read my last post? we can just for get the whole thing. OK? i am a little bit younger and didnt understand it clear ok? i dont want to get even more mad. just please stop. ok?

OpenStudy (mathmale):

Two people have given you the proper formula for the area of a triangle. You are also given the base and the height. Substitute these values into the formula A = (1/2) (base) (height) and you'll have the result you wanted.

OpenStudy (mathmale):

I have no idea what you mean by "parentheses in the end." Never mind. Please, SnowFlake, calculate this area and move on. I'm glad you want to be friends. So do I.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thank you 4 all ur help

OpenStudy (mathmale):

My pleasure. Really.

OpenStudy (anonymous):

omigod i just realized something you told me i understod i looked back in the comments and relizen i wasnt deviding it by two. thank you @mathmale

OpenStudy (mathmale):

Hope you've arrived at a result that seems right to you. Rest assured that I'm glad for you if you have, and that I appreciate your attitude change. :)

OpenStudy (anonymous):

:)